# Unified Primitive Proxies for Structured Shape Completion

### v1

@@ 4.3 Planar Primitive Results
Method PolyFit [nan2017polyfit] KSR [bauchet2020kinetic] COMPOD [sulzer2024concise]
@@
AdaPoinTr [yu2023adapointr] 4.87 10.61 0.934 0.85 4.50 9.56 0.939 0.07 5.92 18.94 0.917 0.15
ODGNet [cai2024orthogonal] 3.97 9.09 0.947 0.87 4.41 9.88 0.947 0.61 4.71 13.81 0.940 1.77
SymmComplete [yan2025symmcompletion] 7.55 14.74 0.893 24.98 13.09 24.45 0.803 38.94 9.73 22.54 0.878 20.48
PaCo [chen2025parametric] 4.89 10.74 0.932 0.54 4.47 10.20 0.934 0.17 4.71 11.73 0.932 0.00
@@ Acknowledgments
This work was supported by TUM Georg Nemetschek Institute under the AI4TWINNING project. We thank Liangliang Nan for helpful discussions.
@@ B Implementation Details
The point pathway follows the AdaPoinTr backbone [yu2023adapointr] with its default depth and hyperparameters. Input points are grouped into local neighborhoods, encoded with self-attention, and decoded from learned point queries. The decoder produces $U=512$ shape features, which a lightweight reconstruction head expands into 8,192 completed points. As in AdaPoinTr, we use denoising queries during training for an auxiliary denoising loss and drop them at inference. The primitive pathway consumes the same $U$ shape features. We use $K=40$ learnable primitive queries, contextualized by a 4-layer Transformer decoder with 8 attention heads and a hidden size of 128. For each query, a prediction head outputs a primitive type, a soft mask over the 512 shape features, and 10 coefficients of a homogeneous quadric.
@@
$\alpha_{1}=\begin{cases}0.05,&\text{if }c_{i}\neq\emptyset,\\ 0.01,&\text{otherwise.}\end{cases}$ (S1)
where $c_{i}$ denotes the ground-truth primitive type. On ABC-multi, shapes contain on average about 8 primitives while we use $K=40$ proxies, so the ratio for no-object vs. valid types roughly matches the expected proportion and keeps their aggregate loss contributions comparable. The impact of key hyperparameters is analyzed in Sec. C.
@@ C.1 More Ablations
We vary the number of primitive proxies $K\in\{30,40,50\}$ and the no-object weight $\alpha_{1}(c_{i}=\emptyset)$ while keeping all other settings fixed. As reported in Tab. S1, $K=40$ with a no-object weight of 0.01 achieves the best results.
@@
$K$ $\alpha_{1}(c_{i}=\emptyset)$ Params (M) CD $\downarrow$ HD $\downarrow$ NC $\uparrow$ FR $\downarrow$
@@
We assess the effect of the confidence threshold applied to the scores $s_{k}$ in Eq. 14, varying the pruning value in $\{0.3,0.5,0.7\}$ at inference time. As summarized in Tab. S2, performance is fairly stable across thresholds, and $s_{k}\geq 0.5$ gives the best overall results.
Table S2: Effect of confidence threshold. Changing the pruning cutoff on $s_{k}$ within a reasonable range has little effect on performance, while using $s_{k}\geq 0.5$ gives the best results.
$s_{k}\geq$ CD $\downarrow$ HD $\downarrow$ NC $\uparrow$ FR $\downarrow$
@@
Tab. 4(b) compares UniCo with and without a projection-based post-processing step that projects completed points onto their predicted primitives to enforce stricter planar geometry, with PaCo [chen2025parametric] as a reference. On both ABC-plane and Building-PCC, UniCo without projection already improves over PaCo, and projection brings only modest additional gains. For fairness, the main paper therefore reports UniCo without projection and treats the projected variant as an optional refinement for slightly sharper surfaces. Fig. S1 shows that PaCo tends to under-represent small primitives and fine details, whereas UniCo produces more uniform point distributions and cleaner, well-aligned primitive layouts across scales. Projection mainly sharpens surfaces and does not change this overall qualitative picture.
@@ C.3 Computational Efficiency
Tab. S6 reports per-scan runtimes on ABC-plane. UniCo processes each partial input in 27.6 ms end-to-end, roughly twice as fast as the strongest pointwise competitor, ODGNet [cai2024orthogonal], and faster than the structured competitor PaCo [chen2025parametric]. All other completion methods, except PaCo and UniCo, require an additional primitive extraction stage. To quantify this overhead, we evaluate RANSAC [schnabel2007efficient], HPNet [yan2021hpnet], and PTv3 [wu2024ptv3] on 100 randomly selected ABC-multi samples completed by ODGNet. UniCo is the fastest end-to-end pipeline.
@@
Method Params Latency RANSAC [schnabel2007efficient] HPNet [yan2021hpnet] PTv3 [wu2024ptv3]
@@
AdaPoinTr [yu2023adapointr] 32.5 23.9 32.5 124.9    33.8 1280.9    208.5 169.3
ODGNet [cai2024orthogonal] 11.5 53.6 11.5 154.6    12.8 1310.6    187.5 199.0
SymmComplete [yan2025symmcompletion] 13.3 20.0 13.3 121.0    14.6 1277.0    189.4 165.4
PaCo [chen2025parametric] 41.4 29.8 41.4 29.8  41.4 29.8  41.4 29.8
@@ D Datasets
For ABC-multi, we randomly select 30,000 single-piece watertight CAD models from the ABC dataset [koch2019abc], covering a broad spectrum of primitive configurations, from simple shapes with only a few primitives to complex assemblies with several tens of parts. Planes and cylinders dominate the primitive inventory, while cones and spheres provide additional geometric variety. This yields a diverse, structured benchmark for primitive-based completion. A quantitative summary is provided in Tab. S7.
@@
In addition, we use ABC-plane [chen2025parametric], a plane-only subset of ABC, and Building-PCC [gao2024building], an airborne LiDAR dataset of roughly 50,000 urban buildings with noise and occlusions. These datasets complement ABC-multi by providing plane-only CAD assemblies and real-world scans for evaluation.
@@ E Metrics
For failed reconstructions, we follow the established protocol [chen2025parametric] and evaluate against the unit cube. Below we detail the additional primitive-level metrics used in Tab. 3, following established practice [guo2022complexgen, liu2024point2cad, li2019supervised]. We denote a matched primitive pair by $(k^{*},g^{*})\in\mathcal{M}$:

### v2

@@ 4.3 Planar Primitive Results
Method PolyFit [38] KSR [1] COMPOD [45]
@@
AdaPoinTr [70] 4.87 10.61 0.934 0.85 4.50 9.56 0.939 0.07 5.92 18.94 0.917 0.15
ODGNet [3] 3.97 9.09 0.947 0.87 4.41 9.88 0.947 0.61 4.71 13.81 0.940 1.77
SymmComplete [63] 7.55 14.74 0.893 24.98 13.09 24.45 0.803 38.94 9.73 22.54 0.878 20.48
PaCo [7] 4.89 10.74 0.932 0.54 4.47 10.20 0.934 0.17 4.71 11.73 0.932 0.00
@@ Acknowledgments
This work was supported by TUM Georg Nemetschek Institute under the AI4TWINNING project. We thank Jing-En Jiang for support with PrimFit and Liangliang Nan for helpful discussions.
@@ B Implementation Details
The point pathway follows the AdaPoinTr backbone [70] with its default depth and hyperparameters. Input points are grouped into local neighborhoods, encoded with self-attention, and decoded from learned point queries. The decoder produces $U=512$ shape features, which a lightweight reconstruction head expands into 8,192 completed points. As in AdaPoinTr, we use denoising queries during training for an auxiliary denoising loss and drop them at inference. The primitive pathway consumes the same $U$ shape features. We use $K=40$ learnable primitive queries, contextualized by a 4-layer Transformer decoder with 8 attention heads and a hidden size of 128. For each query, a prediction head outputs a primitive type, a soft mask over the 512 shape features, and 10 coefficients of a homogeneous quadric.
@@
$\alpha_{1}=\begin{cases}0.05,&\text{if }c_{g}\neq\emptyset,\\ 0.01,&\text{otherwise.}\end{cases}$ (S1)
where $c_{g}$ denotes the ground-truth primitive type. On ABC-multi, shapes contain on average about 8 primitives while we use $K=40$ proxies, so the ratio for no-object vs. valid types roughly matches the expected proportion and keeps their aggregate loss contributions comparable. The impact of key hyperparameters is analyzed in Sec. C.
@@ C.1 More Ablations
We vary the number of primitive proxies $K\in\{30,40,50\}$ and the no-object weight $\alpha_{1}(c_{g}=\emptyset)$ while keeping all other settings fixed. As reported in Tab. S1, $K=40$ with a no-object weight of 0.01 achieves the best results.
@@
$K$ $\alpha_{1}(c_{g}=\emptyset)$ Params (M) CD $\downarrow$ HD $\downarrow$ NC $\uparrow$ FR $\downarrow$
@@
We assess the effect of the confidence threshold applied to the scores $s_{k}$ in Eq. 14, varying the pruning value in $\{0.3,0.5,0.7\}$ at inference time. As summarized in Tab. S2, performance is fairly stable across thresholds, and $s_{k}>0.5$ gives the best overall results.
Table S2: Effect of confidence threshold. Changing the pruning cutoff on $s_{k}$ within a reasonable range has little effect on performance, while using $s_{k}>0.5$ gives the best results.
$s_{k}>$ CD $\downarrow$ HD $\downarrow$ NC $\uparrow$ FR $\downarrow$
@@
Tab. 4(b) compares UniCo with and without a projection-based post-processing step that projects completed points onto their predicted primitives to enforce stricter planar geometry, with PaCo [7] as a reference. On both ABC-plane and Building-PCC, UniCo without projection already improves over PaCo, and projection brings only modest additional gains. For fairness, the main paper therefore reports UniCo without projection and treats the projected variant as an optional refinement for slightly sharper surfaces. Fig. S1 shows that PaCo tends to under-represent small primitives and fine details, whereas UniCo produces more uniform point distributions and cleaner, well-aligned primitive layouts across scales. Projection mainly sharpens surfaces and does not change this overall qualitative picture.
@@
Robustness across backbones.
We additionally swap the point completion backbone $f_{\mathrm{point}}$ to PoinTr [69] or ODGNet [3]. UniCo retains consistent gains over the corresponding completion+extractor baselines. This shows that the improvements are driven by UniCo’s primitive pathway and persist across backbones.
Table S5: Robustness across backbones. UniCo retains consistent gains over the corresponding completion+extractor baselines.
Method $f_{\mathrm{point}}=\text{PoinTr}$ [69] $f_{\mathrm{point}}=\text{ODGNet}$[3]
[…]
Compl.+Extr. 6.58 0.791 11.27 4.33 0.873 7.41
[…]
@@ C.2 Transferability
Moreover, UniCo generalizes to unseen freeform geometries in real OmniObject3D scans [58] without training on them, by compactly approximating them with the available primitives, as shown in Fig. S2.
Figure S2: OOD generalization via approximation. UniCo trained on synthetic data with a fixed primitive set generalizes to real-world point clouds with unseen geometry by approximation.
@@ C.3 Computational Efficiency
Tab. S7 reports per-scan runtimes on ABC-plane. UniCo processes each partial input in 27.6 ms end-to-end, roughly twice as fast as the strongest pointwise competitor, ODGNet [3], and faster than the structured competitor PaCo [7]. All other completion methods, except PaCo and UniCo, require an additional primitive extraction stage. To quantify this overhead, we evaluate RANSAC [42], HPNet [64], and PTv3 [59] on 100 randomly selected ABC-multi samples completed by ODGNet. UniCo is the fastest end-to-end pipeline.
@@
Method Params Latency RANSAC [42] HPNet [64] PTv3 [59]
@@
AdaPoinTr [70] 32.5 23.9 32.5 124.9    33.8 1280.9    208.5 169.3
ODGNet [3] 11.5 53.6 11.5 154.6    12.8 1310.6    187.5 199.0
SymmComplete [63] 13.3 20.0 13.3 121.0    14.6 1277.0    189.4 165.4
PaCo [7] 41.4 29.8 41.4 29.8  41.4 29.8  41.4 29.8
@@
C.4 Failure cases
Typical failure modes involve missing small primitives and misclassified semantics, with examples shown in Fig. S3. These cases are overall rare and occur less often than in competing methods (see Tab. 3).
Figure S3: Failure cases. Typical failure cases involve missing small primitives and misclassified semantics.
@@ D Datasets
For ABC-multi, we randomly select 30,000 single-piece watertight CAD models from the ABC dataset [24], covering a broad spectrum of primitive configurations, from simple shapes with only a few primitives to complex assemblies with several tens of parts. Planes and cylinders dominate the primitive inventory, while cones and spheres provide additional geometric variety. This yields a diverse, structured benchmark for primitive-based completion. A quantitative summary is provided in Tab. S8.
@@
In addition, we use ABC-plane [7], a plane-only subset of ABC, and Building-PCC [15], an airborne LiDAR dataset of roughly 50,000 urban buildings with noise and occlusions. These datasets complement ABC-multi by providing plane-only CAD assemblies and real-world scans for evaluation.
@@ E Metrics
For failed reconstructions, we follow the established protocol [7] and evaluate against the unit cube. Below we detail the additional primitive-level metrics used in Tab. 3, following established practice [16, 36, 30]. We denote a matched primitive pair by $(k^{*},g^{*})\in\mathcal{M}$: